# High-energy Emission from Turbulent Electron-ion Coronae of Accreting Black Holes

### v1

@@ IV.2 Energy Partitioning, Nonthermal Acceleration and Emission Spectrum
Figure 4 also shows that higher magnetizations lead to more pronounced nonthermal features. With the exception of the low-compactness simulation ($\ell\approx 0.6$), the nonthermal ions are accelerated up to the maximum energy set by the size of the simulation box $E_{\rm max}=m_{\rm i}c^{2}\sigma_{\rm i}^{1/2}s_{\rm esc}/d_{\rm i}$ (dotted vertical lines in the middle panel of Fig. 4), where the particle gyroradius $\rho_{\max}=s_{\rm esc}$. The local power-law index in the nonthermal parts of the ion distribution is $p\gtrsim 3$. The electron power-law index hardens with growing magnetization (or compactness). For $\ell>1$, we find $p\approx 3.5$ for electrons. This is broadly consistent with our theoretical discussion from Sec. II.3, where we argue that electrons are predominantly accelerated at intense reconnecting current sheets, in which case a power-law index not far from $p\approx 3.5$ is expected for $\sigma_{\rm i}\sim 0.1$ and $\delta B/B_{0}\sim 1$ (Comisso, 2024). It can be also noticed that in between the electron thermal peak and the power-law tail lies a suprathermal range, where the spectrum is relatively flat. This could be in part related to the hard power-law range expected for high-$\sigma_{\rm e}$ reconnection (see Sec. II.3), but in a turbulent volume the hard power law could be obscured by contributions from particles outside the current sheets, where the spectrum is softer. For a more detailed analysis of particle injection into nonthermal populations near shocks and current sheets see Appendix A.
@@
Figure 5: Comparison of predicted emission spectra with X-ray observations of NGC 4151. Dashed curves show the intrinsic steady-state spectra obtained directly from our PIC simulations with different strengths of the radiative compactness $\ell\approx 0.6,\,2.5,\,8.5$. Solid curves show the spectra corrected for absorption and reflection. Shaded bands indicate the ranges over which the simulated spectra vary during the averaging interval from $t\approx 4S/v_{\rm A}$ to $t\approx 10S/v_{\rm A}$.
Our simulated spectra are in excellent agreement with observations of NGC 4151. This demonstrates that our kinetic simulations probe the observationally relevant regime of turbulence, and so the obtained physical insight (e.g., concerning the energy partitioning) can be put into the context of real sources. The intrinsic X-ray photon index is close to $\Gamma_{\rm x}=1.77$ (dotted black line in Fig. 5), which is in reasonable agreement with previous studies of NGC 4151 (e.g., by Lubiński et al., 2016, who found $\Gamma_{\rm x}=1.76$). For our choice of model parameters (see Sec. III) and for the measured values of $q_{\rm i}$, Eq. (10) predicts $\Gamma_{\rm x}\approx 1.76,\,1.77,\,1.75$ in runs with $\sigma_{\rm i}=0.035,\,0.1,\,0.19$, respectively. The peak of the (intrinsic) spectrum is near 80 keV, which is as well in agreement with our theoretical estimate $E_{\rm peak}\sim m_{\rm e}c^{2}\left\{[1+3/(4\tau_{\rm T}(1\!\!\;+\!\!\;\tau_{\rm T}))]^{1/2}-1\right\}\sim 90$ keV. Thus, our numerical results are well aligned with theoretical expectations for the model.
Figure 5 also shows that the predicted X-ray spectrum is degenerate with respect to model parameters, in the sense that simulations employing different values of $\ell$ can be matched to observations. However, the differences become apparent above the peak of the spectrum, where larger values of the radiative compactness lead to a more pronounced MeV tail. As discussed in Sec. II.3, the MeV tail is shaped by nonthermal electrons, which are sensitive to the microphysical details of the dissipation. In particular, larger values of the compactness imply higher electron magnetizations, which give rise to harder nonthermal tails from electrons accelerated at reconnecting current sheets. Moreover, in the range around 1 MeV the spectrum is nearly unaffected by the (parameter-dependent) reflection modeling. Thus, future measurements with MeV-band instruments (Tomsick et al., 2024) can provide important physical constraints on the emission mechanism. Past attempts to measure the MeV tail in NGC 4151 (Zdziarski et al., 1996) suggest that the spectrum indeed extends into the MeV range, in broad agreement with our predictions for $\ell>1$.
@@ V Discussion and Conclusions
We developed a local emission model of black-hole coronae based on dissipation in a strongly turbulent electron-ion plasma. Our theoretical estimates (Sec. II) are informed by a set of radiative PIC simulations (Sec. IV), which include self-consistent Compton scattering, as well as injection and diffusive escape of charged particles and photons. The simulations allow us to track from first principles the turbulent partitioning of energy among ions, electrons, and photons. We show that the turbulent cascade regulates itself into a trans-sonic state, which features sporadic formation of magnetized shocks (Sec. IV.1). In steady state, the ion kinetic temperature far exceeds the electron kinetic temperature ($T_{\rm i}\gg T_{\rm e}$). Complementary to alternative arguments, based e.g. on assuming that the ions are near the virial temperature (e.g., Bambic et al., 2024), we interpret the two-temperature state as a natural outcome of the interplay between the dissipative microphysics and radiation. In order to maintain the expected steady-state ion kinetic temperature, an order-unity fraction of the available power needs to be deposited into the ions (see Sec. II.2). In our simulations, we measure $0.6\lesssim q_{\rm i}\lesssim 0.7$ for the ion heating fraction, which shows that most of the dissipated power is deposited into the ions.
@@
While our model gives an apparent good match to the observed X-ray spectrum of a bright AGN, it falls short of being able to predict the emission in lower-energy bands, such as the UV/optical (e.g., Steffen et al., 2006; Duras et al., 2020) or the mm band (e.g., Kawamuro et al., 2022; Ricci et al., 2023), and their connection to the X-ray luminosity. We stress that our local model only tracks the soft photons intercepted by the hot corona. Realistically, only a fraction of the UV/optical photons from the disk is actually intercepted by the corona (Done et al., 2007). Similarly, the recently observed mm emission may come from a large-scale outflow launched from the corona (Hankla et al., 2025). Thus, accurate multi-wavelength predictions require a global model for the extended accretion disk and its corona.
Finally, we find that a strongly turbulent corona can efficiently accelerate nonthermal particles (Sec. IV.2). We investigate the origins of nonthermal acceleration and show that the electrons are preferentially injected into nonthermal power laws at intense current sheets, while the ions can be injected at both shocks and current sheets (Appendix A). In our limited-size simulation boxes, the nonthermal ions can be accelerated up to the system size (Hillas) limit, albeit with a relatively steep spectrum (with power-law index $p\gtrsim 3$). Simulations with significantly larger boxes are required to confidently predict the shape of the nonthermal ion (i.e., proton) spectrum up to the neutrino-producing TeV range. However, regardless of the exact shape of the ion spectrum, the maximum fraction of the coronal power deposited into cosmic rays is limited from above by the ion heating fraction $q_{\rm i}$. The relatively large values of the measured heating fraction ($0.6\lesssim q_{\rm i}\lesssim 0.7$) present a quite favorable situation for producing energetically significant amounts of cosmic-ray particles.
@@
We gratefully acknowledge helpful discussions with F. Bacchini, L. Sironi, R. Mbarek, B. Ripperda, H. Krawczynski, A. Hankla, E. Gorbunov, D. Caprioli, L. Comisso, and A. Levinson. D.G. is supported by the Research Foundation–Flanders (FWO) Senior Postdoctoral Fellowship 12B1424N. This work was supported by NASA (grant 80NSSC22K1054) and the Simons Foundation (grant 00001470; A.P.), and was facilitated by the Multimessenger Plasma Physics Center (MPPC; A.P. and A.B.) under NSF grant No. PHY-2206607. A.P. additionally acknowledges support from an Alfred P. Sloan Fellowship and a Packard Foundation Fellowship in Science and Engineering. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the FWO and the Flemish Government. We acknowledge LUMI–BE for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium, through a LUMI–BE Regular Access call. Simulations were additionally performed on NASA Pleiades (GID s2056).
@@ Appendix A Analysis of Shocks and Current Sheets
Figure 7: Injection of particles into nonthermal populations near shocks and current sheets. On the top left we show the out-of-plane electric current $J_{z}$ and the divergence of the bulk velocity $\nabla\cdot{\boldsymbol{\beta}}_{\rm bulk}$ in a fraction of the 2D simulation domain. On the bottom left we show the density of nonthermal electrons $n_{\rm e,nt}$ and nonthermal ions $n_{\rm i,nt}$. The red and green squares mark the representative locations of a current sheet and a shock, respectively, from which we extract the local particle distributions shown on the bottom right side of the figure. On the top right side we show the mean ion and electron pitch angle relative to $\mathbf{B}$, computed in the local $E\times B$ frame.
@@ Appendix B Photon-photon and Ion-electron Collision Rates
Rapid Coulomb collisions between ions and electrons can lead to equilibration of their temperatures. It is therefore useful to check when collisions are sufficiently infrequent to allow for the development of a two-temperature corona with $T_{\rm i}\gg T_{\rm e}$. For simplicity, we employ analytical estimates for a Maxwellian plasma with dimensionless species temperatures $\theta_{\rm s}=T_{\rm s}/(m_{\rm s}c^{2})$ (Stepney, 1983; Stepney & Guilbert, 1983), where in practice $\theta_{\rm s}$ is to be understood as an effective kinetic temperature of the (nonthermal) particle distribution. The collisional ion-electron relaxation time scale, on which hot ions lose energy to cool electrons, can be obtained as $t_{\rm ie}\simeq\left|\kappa_{\rm e}n_{0}(T_{\rm i}-T_{\rm e})/({\rm d}U_{\rm e}/{\rm d}t)\right|$ (Spitzer, 1962), where ${\rm d}U_{\rm e}/{\rm d}t$ is the collisional electron energy exchange rate per unit volume, and $\kappa_{\rm e}=1/(\Gamma_{\rm e}-1)$. For non-relativistic or very mildly relativistic particles we can approximate (Stepney, 1983):
@@ Appendix C Turbulent Mode Decomposition
where $\hat{\mathbf{k}}_{\perp}=\mathbf{k}_{\perp}/k_{\perp}$, $\boldsymbol{\upsilon}=[\Gamma_{\rm bulk}^{3}w_{0}/(\Gamma_{\rm bulk}+1)]^{1/2}{\boldsymbol{\beta}}_{\rm bulk}$ is the weighted fluid velocity, $\Gamma_{\rm bulk}=[1-\beta_{\rm bulk}^{2}]^{-1/2}$ is the fluid Lorentz factor, and $w_{0}$ is the mean proper enthalpy density. [^8] For completeness, we retain the electric field contribution and we use a relativistic definition of the kinetic energy density, but we note that relativistic effects are mild in the regime of interest. The last term in Eq. (C4) can be written in real space as $\delta n\delta p/(2n_{0})$. It can be identified as the energy density of acoustic oscillations, which naturally enter the total energy definition when the cascade is compressible (e.g., see Schekochihin et al., 2009).

### v2

@@ IV.2 Energy Partitioning, Nonthermal Acceleration and Emission Spectrum
Figure 4 also shows that higher magnetizations lead to more pronounced nonthermal features. Nonthermal ions with energies greater than 4 times the mean of their distribution contain roughly 1/5 of the total kinetic energy for $\ell\approx 0.6$ and about 1/4 for $\ell\approx 2.5,\,8.5$. With the exception of the low-compactness simulation ($\ell\approx 0.6$), the nonthermal ions are accelerated up to the maximum energy set by the size of the simulation box $E_{\rm max}=m_{\rm i}c^{2}\sigma_{\rm i}^{1/2}s_{\rm esc}/d_{\rm i}$ (dotted vertical lines in the middle panel of Fig. 4), where the particle gyroradius $\rho_{\max}=s_{\rm esc}$. The local power-law index in the nonthermal parts of the ion distribution is $p\gtrsim 3$. The electron power-law index hardens with growing magnetization (or compactness). For $\ell>1$, we find $p\approx 3.5$ for electrons. This is broadly consistent with our theoretical discussion from Sec. II.3, where we argue that electrons are predominantly accelerated at intense reconnecting current sheets, in which case a power-law index not far from $p\approx 3.5$ is expected for $\sigma_{\rm i}\sim 0.1$ and $\delta B/B_{0}\sim 1$ (Comisso, 2024). It can be also noticed that in between the electron thermal peak and the power-law tail lies a suprathermal range, where the spectrum is relatively flat. This could be in part related to the hard power-law range expected for high-$\sigma_{\rm e}$ reconnection (see Sec. II.3), but in a turbulent volume the hard power law could be obscured by contributions from particles outside the current sheets, where the spectrum is softer. For a more detailed analysis of particle injection into nonthermal populations near shocks and current sheets see Appendix A.
@@
Figure 5: Comparison of predicted emission spectra with X-ray observations of NGC 4151. Dashed curves show the intrinsic steady-state spectra obtained directly from our PIC simulations. Solid color curves show the PIC spectra corrected for absorption and reflection. For $E\lesssim 100$ keV, the PIC spectra obtained for different values of the compactness $\ell$ lie on top of each other and are thus nearly indistinguishable. Shaded bands indicate the ranges over which the simulated spectra vary during the averaging interval from $t\approx 4S/v_{\rm A}$ to $t\approx 10S/v_{\rm A}$. The dotted black curve is the nthcomp thermal Comptonization model of xspec, fitted to the X-ray continuum and corrected for absorption and reflection.
Our simulated spectra are in excellent agreement with observations of NGC 4151. This demonstrates that our kinetic simulations probe the observationally relevant regime of turbulence, and so the obtained physical insight (e.g., concerning the energy partitioning) can be put into the context of real sources. The intrinsic X-ray photon index is close to $\Gamma_{\rm x}=1.77$ (dash-dotted black line in Fig. 5), which is in reasonable agreement with previous studies of NGC 4151 (e.g., by Lubiński et al., 2016, who found $\Gamma_{\rm x}=1.76$). For our choice of model parameters (see Sec. III) and for the measured values of $q_{\rm i}$, Eq. (10) predicts $\Gamma_{\rm x}\approx 1.76,\,1.77,\,1.75$ in runs with $\sigma_{\rm i}=0.035,\,0.1,\,0.19$, respectively. The peak of the (intrinsic) spectrum is near 80 keV, which is as well in agreement with our theoretical estimate $E_{\rm peak}\sim m_{\rm e}c^{2}\left\{[1+3/(4\tau_{\rm T}(1\!\!\;+\!\!\;\tau_{\rm T}))]^{1/2}-1\right\}\sim 90$ keV. Thus, our numerical results are well aligned with theoretical expectations for the model.
Figure 5 also shows that the predicted X-ray spectrum is degenerate, in the sense that simulations employing different values of the compactness $\ell$ can be matched to observations. However, the differences become apparent above the peak of the spectrum, where larger values of the radiative compactness lead to a more pronounced MeV tail. As discussed in Sec. II.3, the MeV tail is shaped by nonthermal electrons, which are sensitive to the microphysical details of the dissipation. The MeV tail generated by nonthermal electrons differs from predictions based on thermal Comptonization models, which predict a sharp cutoff above peak (see black dotted curve in Fig. 5). Larger values of the compactness imply higher electron magnetizations, which give rise to harder nonthermal tails from electrons accelerated at reconnecting current sheets. Moreover, in the range around 1 MeV the spectrum is nearly unaffected by the (parameter-dependent) reflection modeling. Thus, future measurements with MeV-band instruments (Tomsick et al., 2024) can provide important physical constraints on the emission mechanism. Past attempts to measure the MeV tail in NGC 4151 (Zdziarski et al., 1996) suggest that the spectrum indeed extends into the MeV range, in broad agreement with our predictions for $\ell>1$.
@@ V Discussion and Conclusions
We developed a local emission model of black-hole coronae based on dissipation in a strongly turbulent electron-ion plasma. Our theoretical estimates (Sec. II) are informed by a set of radiative PIC simulations (Sec. IV), which include self-consistent Compton scattering, as well as injection and diffusive escape of charged particles and photons. The simulations allow us to track from first principles the turbulent partitioning of energy among ions, electrons, and photons. We show that the turbulent cascade regulates itself into a trans-sonic state, which features sporadic formation of magnetized fast-mode shocks (Sec. IV.1). In steady state, the ion kinetic temperature far exceeds the electron kinetic temperature ($T_{\rm i}\gg T_{\rm e}$). Complementary to alternative arguments, based e.g. on assuming that the ions are near the virial temperature (e.g., Bambic et al., 2024), we interpret the two-temperature state as a natural outcome of the interplay between the dissipative microphysics and radiation. In order to maintain the expected steady-state ion kinetic temperature, an order-unity fraction of the available power needs to be deposited into the ions (see Sec. II.2). In our simulations, we measure $0.6\lesssim q_{\rm i}\lesssim 0.7$ for the ion heating fraction, which shows that most of the dissipated power is deposited into the ions.
@@
While our model gives an apparent good match to the observed X-ray spectrum of a bright AGN, it falls short of being able to predict the emission in lower-energy bands, such as the UV/optical (e.g., Steffen et al., 2006; Duras et al., 2020) or the mm band (e.g., Kawamuro et al., 2022; Ricci et al., 2023), and their connection to the X-ray luminosity. We stress that our local model only tracks the soft photons intercepted by the hot corona. Realistically, only a fraction of the UV/optical photons from the disk is actually intercepted by the corona (Done et al., 2007). Similarly, the recently observed mm emission may come from a large-scale outflow launched from the corona (Hankla et al., 2026). Thus, accurate multi-wavelength predictions require a global model for the extended accretion disk and its corona (e.g., Nagele et al., 2026).
Finally, we find that a strongly turbulent corona can efficiently accelerate nonthermal particles (Sec. IV.2). We investigate the origins of nonthermal acceleration and show that the electrons are preferentially injected into nonthermal power laws at intense current sheets, while the ions can be injected at both shocks and current sheets (Appendix A). In our limited-size simulation boxes, the nonthermal ions can be accelerated up to the system size (Hillas) limit, albeit with a relatively steep spectrum (with power-law index $p\gtrsim 3$). Simulations with significantly larger boxes are required to confidently predict the shape of the nonthermal ion (i.e., proton) spectrum up to the neutrino-producing TeV range. However, regardless of the exact shape of the ion spectrum, our simulations establish that a significant fraction of the coronal power can be converted into cosmic rays. In particular, we find that ions in the nonthermal tail of the distribution amount to roughly $q_{\rm i}/4\sim 1/6$ of the total dissipated power.
@@
We gratefully acknowledge helpful discussions with F. Bacchini, L. Sironi, R. Mbarek, B. Ripperda, H. Krawczynski, A. Hankla, E. Gorbunov, D. Caprioli, L. Comisso, A. Levinson, M. Nhat Ly, M. Lemoine, and C. Nagele. D.G. is supported by the Research Foundation–Flanders (FWO) Senior Postdoctoral Fellowship 12B1424N. This work was supported by NASA (grant 80NSSC22K1054) and the Simons Foundation (grant 00001470; A.P.), and was facilitated by the Multimessenger Plasma Physics Center (MPPC; A.P. and A.B.) under NSF grant No. PHY-2206607. A.P. additionally acknowledges support from an Alfred P. Sloan Fellowship and a Packard Foundation Fellowship in Science and Engineering. A.M.B. is supported by NASA ATP grant 80NSSC24K1229, NSF grant AST-2408199, and Simons Foundation award 446228. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the FWO and the Flemish Government. We acknowledge LUMI–BE for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium, through a LUMI–BE Regular Access call. Simulations were additionally performed on NASA Pleiades (GID s2056).
@@ Appendix A Analysis of Shocks and Current Sheets
Figure 7: Injection of particles into nonthermal populations near shocks and current sheets. On the top left we show the out-of-plane electric current $J_{z}$ and the divergence of the bulk velocity $\nabla\cdot{\boldsymbol{\beta}}_{\rm bulk}$ in a fraction of the 2D simulation domain. On the bottom left we show the density of nonthermal electrons $n_{\rm e,nt}$ and nonthermal ions $n_{\rm i,nt}$. The red and green squares mark the representative locations of a current sheet and a shock, respectively, from which we extract the local particle distributions shown on the bottom right side of the figure. On the top right side we show the mean cosine of the ion and electron pitch angle relative to $\mathbf{B}$, computed in the local $E\times B$ frame.
@@
Composite electron spectrum
In Fig. 8 we show electron distributions extracted from spatial cells where the local value of $T_{\rm e}$ falls into a given range. This further clarifies how the total spectrum is shaped by turbulent fluid motions of moderately hot electrons in the bulk of the volume, and by nonthermal particles at spatially localized hotspots.
Figure 8: Electron energy distributions extracted from locations with different values of the local temperature. Each color curve represents the distribution of particles from spatial cells where $\log_{10}(T_{\rm e})$ falls into a given range (as indicated by the colorbar). The distributions are normalized relative to the number of particles contained in a given band, such that their sum gives the total (box-averaged) distribution (black dashed curve).
As shown in Fig. 8, the most numerous particle population (cyan curve) has a typical proper kinetic temperature of $T_{\rm e}\approx 10^{-1.3}m_{\rm e}c^{2}\approx 0.05\,m_{\rm e}c^{2}$. Such temperature is too low to fully account for the peak of the box-averaged spectrum near $E\approx 0.15m_{\rm e}\,c^{2}$. The discrepancy can be explained by contributions from bulk turbulent motions. Indeed, even in the lowest temperature band (dark blue curve) with $T_{\rm e}\approx 10^{-1.7}m_{\rm e}c^{2}\approx 0.02\,m_{\rm e}c^{2}$ the peak of the distribution is near $E\approx 0.1m_{\rm e}c^{2}$, which indicates that bulk motions put an effective floor on the minimum particle energy. In other words, in the simulation frame the electron energies are boosted by the bulk motions, and therefore the distributions peak at larger values than expected for a given temperature.
At the most extreme hotspots with $T_{\rm e}\approx 10^{-0.3}m_{\rm e}c^{2}\approx 0.5\,m_{\rm e}c^{2}$ (dark red curve in Fig. 8), the electron distribution exhibits a hard power law for $0.1\,m_{\rm e}c^{2}\lesssim E\lesssim 2m_{\rm e}c^{2}$, and a steeper spectrum above this range. The hard power-law range (with $p\lesssim 1$) is reminiscent of spectra obtained in high-$\sigma_{\rm e}$ reconnection studies (Guo et al., 2024; Sironi et al., 2025). Together with the results shown in Fig. 7, we can conclude that turbulent current sheets can produce spectra broadly similar to those expected from dedicated reconnection studies (e.g., Comisso, 2024). However, because the most intense energization sites occupy a small volume, the composite (box-averaged) electron spectrum is significantly softer and different from the spectrum of an isolated current sheet.
@@ Appendix B Photon-photon and Ion-electron Collision Rates
Rapid Coulomb collisions between ions and electrons can lead to equilibration of their temperatures. It is therefore useful to check when collisions are sufficiently infrequent to allow for the development of a two-temperature corona with $T_{\rm i}\gg T_{\rm e}$. For simplicity, we employ analytical estimates for a Maxwellian plasma with dimensionless species temperatures $\theta_{\rm s}=T_{\rm s}/(m_{\rm s}c^{2})$ (Stepney, 1983; Stepney & Guilbert, 1983), where in practice $\theta_{\rm s}$ is to be understood as an effective kinetic temperature of the (nonthermal) particle distribution. The collisional ion-electron relaxation time scale, on which hot ions lose energy to electrons, can be obtained as $t_{\rm ie}\simeq\left|\kappa_{\rm e}n_{0}(T_{\rm i}-T_{\rm e})/({\rm d}U_{\rm e}/{\rm d}t)\right|$ (Spitzer, 1962), where ${\rm d}U_{\rm e}/{\rm d}t$ is the collisional electron energy exchange rate per unit volume, and $\kappa_{\rm e}=1/(\Gamma_{\rm e}-1)$. For non-relativistic or very mildly relativistic particles we can approximate (Stepney, 1983):
@@ Appendix C Turbulent Mode Decomposition
where $\hat{\mathbf{k}}_{\perp}=\mathbf{k}_{\perp}/k_{\perp}$, $\boldsymbol{\upsilon}=[\Gamma_{\rm bulk}^{3}w_{0}/(\Gamma_{\rm bulk}+1)]^{1/2}{\boldsymbol{\beta}}_{\rm bulk}$ is the weighted fluid velocity, $\Gamma_{\rm bulk}=[1-\beta_{\rm bulk}^{2}]^{-1/2}$ is the fluid Lorentz factor, and $w_{0}$ is the mean proper enthalpy density. [^8] For completeness, we retain the electric field contribution and we use a relativistic definition of the bulk kinetic energy density, but we note that relativistic effects are mild in the regime of interest. Note that $|{\boldsymbol{\upsilon}}|^{2}=\Gamma_{\rm bulk}w_{0}(\Gamma_{\rm bulk}-1)$, which reduces to $n_{0}m_{\rm i}c^{2}\beta_{\rm bulk}^{2}/2$ in the non-relativistic limit. The last term in Eq. (C4) can be written in real space as $\delta n\delta p/(2n_{0})$. It can be identified as the energy density of acoustic oscillations, which naturally enter the total energy definition when the cascade is compressible (e.g., see Schekochihin et al., 2009).
@@
For reference, we demonstrate in Fig. 9 how the total energy spectrum (as shown in Fig. 3 of Sec. IV.1) is partitioned among the different forms of energy that appear in Eq. (C4). As expected, at MHD scales ($k_{\perp}\lesssim 1/d_{\rm i}$) most of the energy is contained in the magnetic (red curve) and bulk kinetic energy density (purple curve). At sub-ion scales, the energy is primarily contained in the magnetic and density fluctuations (yellow curve), consistent with expectations for kinetic Alfvén turbulence (e.g., Grošelj et al., 2019). The flattening of the kinetic energy density spectrum at $k_{\perp}d_{\rm i}\sim 10$ can be attributed to PIC noise artifacts.
Figure 9: Total turbulent energy spectrum from Fig. 3 decomposed among the different forms of energy that appear in Eq. (C4).